1/y + 1/y+2
divided by
1/y- 1/y+2

answer:y+1 i think my text book is wrong anyone know if this is the right answer and if it is can u give a full break down on why the answer is y+1

Answer by stanbon(75887)

1/y + 1/(y+2)
divided by
1/y - 1/(y+2)
--------------------------
Simplify the numerator:
= (y+2 + y)/(y(y+2)) = (2y+2)/(y(y+2))
---------------------------------------------
Simplify the denominator:
= (y+2-y)/(y(y+2))
= 2/(y(y+2))
--------------------------
Invert the denominator and multiply to get:
[(2y+2)/(y(y+2))/ * [y(y+2))/2]
---
Cancel the y(y+2) which is common to numerator and denominator:
---
= (2y+2)/2
---
Cancel the common factor of "2" to get:
= 1/(y+1)
=======================
Cheers,
Stan H.